an ammeter and copper voltmeter are connected in series in an electric circuitthrough which a constant direct current flows.the ammeter shows o.525 ampere.if 0.6354 g of Cu is deposited in one hour,what is percentage error of ammeter

Grade:12

1 Answers

Vikas TU
7 years ago
For Cu to be deposited it would have gone under the oxidation to Cu2+ and the n-factor therfore here would be 2.
From the Farday's electrochemistry law,
(0.6354/63.5)*2 = i*60*60/96500
i = 0.536 A
prcentage error => (0.536-0.525)*100/0.536 = 2.05%